x+(2x/(x-2))=((3x-2)/(x-2))
Answer by stanbon(75887)

x+(2x/(x-2))=((3x-2)/(x-2))
---------
Multiply thru by x-2 to get:
x(x-2) + 2x = 3x-2
---------
x^2 -2x + 2x = 3x-2
-------
x^2-3x+2 = 0
Factor:
(x-2)(x-1) = 0
----
Possible answers:: x = 2 or x = 1
---
Note x = 2 cannot be an answer as that would create zero denominators
in the original problem.
----
Ans: x = 1 is the only solution.
===================================
Cheers,
Stan H.
================